# D-LIO: 6DoF Direct LiDAR-Inertial Odometry based on Simultaneous Truncated Distance Field Mapping

### v1

@@ Abstract
This paper presents a new approach for 6DoF Direct LiDAR-Inertial Odometry (D-LIO) based on the simultaneous mapping of truncated distance fields on CPU. Such continuous representation (in the vicinity of the points) enables working with raw 3D LiDAR data online, avoiding the need of LiDAR feature selection and tracking, simplifying the odometry pipeline and easily generalizing to many scenarios. The method is based on the proposed Fast Truncated Distance Field (Fast-TDF) method as a convenient tool to represent the environment. Such representation enables i) solving the LiDAR point-cloud registration as a nonlinear optimization process without the need of selecting/tracking LiDAR features in the input data, ii) simultaneously producing an accurate truncated distance field map of the environment, and iii) updating such map at constant time independently of its size. The approach is tested using open datasets, aerial and ground. It is also benchmarked against other state-of-the-art odometry approaches, demonstrating the same or better level of accuracy with the added value of an online-generated TDF representation of the environment, that can be used for other robotics tasks as planning or collision avoidance. The source code is publicly available at https://anonymous.4open.science/r/D-LIO
@@ I Introduction
This paper presents, to the best of our knowledge, the first 6DoF LiDAR Inertial SLAM method based on simultaneous mapping of truncated distance field in large volumes. Although the approach has potential for parallelization, it is able to map the environment and estimate the odometry approaching real time using just CPU. Unlike learning-based methods, the proposed Fast-TDF approach provides an accurate representation of the environment from the first scan.
The paper is structured as follows. First, Section II analyses the current state of the art on LiDAR odometry and distance field computation. Later, the proposed Fast Truncated Distance Field method is presented in Section III. Then, Section IV details the LiDAR-Inertial odometry approach based on simultaneous TDF mapping. The proposed methods are tested and benchmarked using public datasets and open-sourced LiDAR odometry approaches in Section V. Finally, Section VI presents the conclusions and future work.
@@ II-A LiDAR-based Odometry
Modern LiDAR-based odometry relies on high-resolution 3D point clouds to estimate motion, offering a key advantage over cameras by operating independently of ambient lighting conditions. Traditional methods, such as ICP [15] and NDT [16] register consecutive scans by minimizing alignment errors. These have been enhanced by approaches like GICP [8], CT-ICP [17] and KISS-ICP [3] that improve robustness and computational efficiency. Additionally, scan-to-map strategies, including IMLS-SLAM [18] and DLO [5], leverage accumulated map information to enhance localization stability.
To reduce computational cost, feature-based methods extract geometric structures rather than using entire point clouds. LOAM [4] pioneered this approach by detecting edges and planar features, later extended by LeGO-LOAM [6] and by SuMa++ [19], which integrate segmentation and semantic constraints. Other methods, such as F-LOAM [7], further optimize the selection and representation of features for robust odometry.
@@
Despite its advantages, LiDAR-only odometry faces several limitations. In featureless environments, such as tunnels, point cloud registration becomes unreliable due to insufficient constraints. Additionally, LiDAR’s low frame rate (10–20 Hz) makes it less effective for capturing rapid motion dynamics. To mitigate these issues, sensor fusion with complementary modalities is employed to enhance robustness and computational efficiency. LiDAR-inertial odometry is an approach commonly used as IMU provide high-frequency inertial measurements capturing short-term motion dynamics with minimal latency.
Sensor fusion methods can be categorized depending on how each sensor contributes to the estimation of the pose. In loosely coupled approaches, LiDAR and IMU measurements are processed independently, and their outputs are fused at a later stage. These methods offer modularity and robustness but may not fully exploit the complementary nature of both sensors. Examples include LOAM [4] or Lego-LOAM [6].
On the other hand, tightly coupled approaches directly integrate raw IMU and LiDAR measurements within the same optimization framework, enabling continuous state estimation and higher accuracy. Methods such as LIO-SAM [2] or Faster-LIO [23] perform well in feature-sparse environments, where IMU constraints help maintain reliable pose estimates.
Despite advances in LiDAR-based odometry, several challenges still hinder its performance. One of the main issues is the high computational cost of processing dense point clouds, making it difficult to execute registration and optimization algorithms efficiently on hardware-constrained platforms. Another significant challenge is the difficulty of extracting discriminative features in homogeneous environments or repetitive geometries, potentially leading to errors in motion estimation. In such scenarios, direct odometry methods, that leverage the full point cloud information without relying on feature extraction, are often considered as an alternative [5] [19]. Furthermore, filtering and data reduction strategies must balance the need to decrease computational load while preserving essential information for localization.
@@ II-B Distance Fields Computation
Volumetric mapping and Simultaneous Localization and Mapping (SLAM) methods have significantly contributed to robot navigation, planning, and localization across diverse environments. These approaches primarily utilize discrete voxel-based structures, hierarchical data representations, probabilistic models, and implicit neural representations.
Voxel-based methods are extensively used due to their simplicity and ease of implementation. Voxblox [24] efficiently generates Euclidean Signed Distance Fields (ESDF) from Truncated Signed Distance Fields (TSDF), establishing a common baseline in volumetric mapping. FIESTA [25] further enhances computational efficiency by employing occupancy-based wavefront propagation instead of TSDF, thereby reducing complexity. Additionally, Voxfield [9] refines accuracy through non-projective TSDF fusion complemented by normal estimation from input point-cloud data. Despite their popularity and CPU efficiency, these voxel-based solutions exhibit inherent limitations such as scalability issues and discretization errors, reducing their effectiveness and precision in real-time SLAM applications.
Hierarchical data structures such as OpenVDB [10] have been proposed to mitigate scalability constraints. Approaches leveraging OpenVDB include VDB-EDT [26], which represents ESDF through efficient distance transforms. While these hierarchical methods improve scalability and memory usage, their discrete nature still limits smooth interpolation capabilities and continuous optimization.
Implicit neural representations have emerged as continuous alternatives to discrete methods, providing dense and differentiable volumetric maps. DeepSDF [27] introduced a neural-network-based implicit model to represent continuous SDFs. Continual Neural Mapping [28] proposed an incremental neural method employing replay buffers, but real-time incremental updates were first effectively demonstrated by iSDF [12]. More recently, HIO-SDF [13] introduced a hierarchical approach combining global voxel structures with local LiDAR data to train neural models online. TNDF-Fusion [29] compresses city-scale LiDAR maps via a Tri-Pyramid TNDF and NeRF-SLAM [30] delivers real-time dense monocular SLAM with hierarchical NeRFs, yet none of them unifies full 6-DoF LiDAR–inertial odometry with on-the-fly distance field mapping on a CPU. Despite these advances, neural methods typically require significant pre-training, heavy computation, and are limited to small-scale SLAM due to memory and inference constraints.
Gaussian Processes (GP) have been explored as probabilistic alternatives for representing distance fields, offering uncertainty estimation capabilities. Previous approaches have used GP-based models for uncertainty-aware mapping [31], but their inference remains computationally intensive due to the inversion of covariance matrices. To address computational constraints, hierarchical representations such as Octrees have been employed [31]. VDB-GPDF [14] leverages a logarithmic Gaussian Processes embedded within the OpenVDB structure, generating continuous ESDFs with uncertainty measures. Nonetheless, the computational complexity associated with GP inference remains a significant obstacle for real-time implementation.
@@
In contrast, the approach proposed in this paper, Direct LiDAR Inertial Odometry (D-LIO), explicitly targets these limitations. D-LIO simplifies the odometry pipeline by eliminating the necessity for explicit LiDAR feature selection and tracking. Moreover, it maintains the environment map update process in constant time, irrespective of map size, significantly enhancing scalability. By directly fusing raw LiDAR point-cloud data into an optimized Fast Truncated Distance Field (Fast-TDF) representation, our method eliminates the additional computational overhead typically required to convert TSDF to ESDF, thus facilitating real-time integration into SLAM systems.
@@ III Fast Truncated Distance Field
Distance Fields (DFs) are able to provide the closest obstacle to a point in space, no matter how far we are from the obstacle. On the other hand, Truncated Distance Fields (TDFs) are only useful when the coordinate has an obstacle under its truncation distance. While this might be seen as a great advantage of DFs for localization, practical aspects limits and even dismiss such advantage because far points are normally rejected by the outlier rejection process during point cloud registration. Thus, if the truncation distance is properly tuned according to the localization approach at hand, TDFs might be a very efficient representation of the distance field.
@@ III-A Binary kernel
As an illustrative example, let’s assume we have three different distance values represented as a bit-set mask of 8 bits, these values are $00111111$ ($L_{1}=6$), $00001111$ ($L_{1}=4$) and $00000001$ ($L_{1}=1$). The shortest distance among these values can be easily computed by computing the bitwise-AND operation without performing value comparison:
@@ III-B TDF grid map
Our TDF grid map will be represented as a 3D matrix of fixed resolution. The size of the grid is determined by the spatial distribution of 3D points. Every grid cell contains the $L_{1}$ distance to the closest point of the 3D map to that cell, or the truncation distance if the closest point is far.
The truncation distance will be defined by three factors: the number of bits used to represent the bit-set mask, the size of the kernel, and the grid resolution. The number of bits of the mask must be equal to the maximum $L_{1}$ distance into the kernel in order to make the approach efficient, so the truncation distance will be finally given by the size of the kernel and the resolution.
III-C Grid map initialisation and access
As previously introduced, the grid map is a 3D matrix containing a binary mask for each cell in the space. The number of bits per cell will be defined by the binary mask size itself. As we have a dense representation of the world, it is not recommended to use sparse representations such as kd-trees [32] or Octomap [33]. Instead, we use a large memory array to represent the space at fixed resolution. A simple hash function is used to transform the 3D coordinates into the corresponding index in the array (see details in the source code). In this way, the time required for grid access is extremely small, and more importantly, constant.
Such grid map must be correctly initialized in order to properly work with the binary kernel bitwise-AND operation. Thus, the whole grid must be initialized to the truncation distance, that is, a mask with all bits set. This value will be updated by the binary kernel as soon as the value to store into the grid cell is smaller than the truncation distance.
@@ III-E Distance interpolation
Following [34], we make use of trilinear interpolation in order to estimate the distance field beyond the grid cell resolution. Trilinear interpolation is used as a trade-off between computation and accuracy. Thus, tricubic interpolation offers better accuracy, however its computation load in each grid cell prevents its online use.
Using trilinear interpolation also offers an easy and quick distance field gradient approximation. Given such gradient approximation, we can analytically compute complex gradients with respect to third-party variables, as robot position or orientation for localization. Contrary to [34], this interpolation is computed online.
@@ IV Direct 3D LiDAR-Inertial Odometry
Figure 2 shows a block diagram of the D-LIO approach. The algorithm incrementally estimates the system pose using an Extended Kalman Filter (EKF) by integrating inertial measurements for pose prediction and point cloud to map registration for pose update. Thus, for each incoming point cloud, the system performs a cloud registration by aligning it with the environment mapped so far based on the Truncated Distance Field (TDF), using the EKF estimate as initial guess. This alignment is used to constraint the EKF’s pose and velocity by means of state update. Whenever the estimated pose variation exceeds a predefined threshold, the TDF map is updated by incorporating new information. This updated map then facilitates a more precise alignment of subsequent point clouds during the registration process.
@@ IV-A Inertial State Integration
Following LiDAR odometry common practices [35][36], inertial sensors can be used to improve state estimation between scans. Inertial integration provides a good initial solution for LiDAR scan alignment, which benefits the subsequence scan-matching process. In addition, high-frequency inertial integration enables LiDAR scan deskewing, which also positively impacts the quality of the scan-matching process. We make use of an EKF with the following state vector:
@@ V Experimental results
The algorithm has been tested on two datasets, the VIRAL Dataset [38] and the Newer College Dataset [39], to evaluate its ability to correctly localize in different environments using diverse trajectories that challenge its performance.
The accurate estimation of odometry has been primarily validated with the VIRAL Dataset, as it features more challenging environments, including the presence of trees, reflective or transparent surfaces, and complex aerial trajectories that test the algorithm’s stability. On the other hand, with the Newer College Dataset, the correct generation of a grid has been assessed in a significantly larger environment in order to demonstrate the algorithm’s scalability concerning map size and number of points. A supplementary video presents qualitative results and 3D reconstructions. [^1] https://youtu.be/iIVCYR85AHg
In both datasets, the Absolute Translation Error (ATE) has been computed by comparing the estimated trajectory with the ground truth. We used the benchmarking tools provided by the VIRAL dataset to ensure a consistent and reliable trajectory evaluation.
The results obtained on both datasets are presented in Table I together with other approaches, such as A-LOAM [40], M-LOAM [41], FAST-LIO2 [36], KISS-ICP [3], and LIO-SAM[2]. The metrics of A-LOAM, M-LOAM, and LIO-SAM were taken from the VIRAL-SLAM paper [42], and those for KISS-ICP from [43]. The LIO-SAM results in quad-easy sequence are taken from NV-LIOM [44]. Please, notice that LIO-SAM performs back-end optimization, which makes the comparison with pure odometry not entirely fair. However, we have included its results to demonstrate that our approach, despite not using back-end optimization, achieves comparable performance. As shown in Table I, our method attains the lowest ATE in the majority of sequences and ranks second-best in the remainder, highlighting its consistent accuracy and robustness across diverse and challenging trajectories.
These experimental results made use of FastTDF configured with a $40\times 40\times 40$ binary kernel using 64 bit mask encoding. Together with a grid resolution of 0.05 m, the truncation distance of this kernel is approximately 3 meters.
V-A VIRAL Dataset
Our algorithm has been tested on the sequences “eee”,“nya”, and “sbs”, each with three sub-trajectories. These sequences cover diverse environments, including indoor and enclosed spaces, open areas with varying building heights, and surfaces that challenge LiDAR odometry, such as glass and semi-transparent materials.
The vehicle employed to record this dataset is a DJI M600 Hexacopter equipped with several sensors, including the IMU “VectorNav VN100” and the horizontal and vertical LiDAR “Ouster OS1-16 gen 1”, with a combined point cloud density of approximately 20 k points per frame. For these experiments, the map size was set to $60\times 60\times 25$ meters, and the parameters $\lambda=1.0$, $t_{\text{th}}=2\,\text{m}$, and $q_{\text{th}}=25^{\circ}$.

### v2

@@ Abstract
This paper presents a new approach for 6DoF Direct LiDAR-Inertial Odometry (D-LIO) based on the simultaneous mapping of truncated distance fields on CPU. Such continuous representation (in the vicinity of the points) enables working with raw 3D LiDAR data online, avoiding the need of LiDAR feature selection and tracking, simplifying the odometry pipeline and easily generalizing to many scenarios. The method is based on the proposed Fast Truncated Distance Field (Fast-TDF) method as a convenient tool to represent the environment. Such representation enables i) solving the LiDAR point-cloud registration as a nonlinear optimization process without the need of selecting/tracking LiDAR features in the input data, ii) simultaneously producing an accurate truncated distance field map of the environment, and iii) updating such map at constant time independently of its size. The approach is tested using open datasets, aerial and ground. It is also benchmarked against other state-of-the-art odometry approaches, demonstrating the same or better level of accuracy with the added value of an online-generated TDF representation of the environment, that can be used for other robotics tasks as planning or collision avoidance. The source code is publicly available at https://github.com/robotics-upo/D-LIO.git
@@ I Introduction
This paper presents, to the best of our knowledge, the first 6DoF LiDAR-Inertial odometry method based on simultaneous mapping of truncated distance field in large volumes. Although the approach has potential for parallelization, it is able to map the environment and estimate the odometry approaching real time using just CPU. Unlike learning-based methods, the proposed Fast-TDF approach provides an accurate representation of the environment from the first scan.
The paper is structured as follows. First, Section II analyzes the current state of the art on LiDAR odometry and distance field computation. Later, the proposed Fast Truncated Distance Field method is presented in Section III. Then, Section IV details the LiDAR-Inertial odometry approach based on simultaneous TDF mapping. The proposed methods are tested and benchmarked using public datasets and open-sourced LiDAR odometry approaches in Section V. Finally, Section VI presents the conclusions and future work.
@@ II-A LiDAR-based Odometry
Modern LiDAR-based odometry relies on high-resolution 3D point clouds to estimate motion, offering a key advantage over cameras by operating independently of ambient lighting conditions. Traditional methods, such as ICP [14] and NDT [15], register consecutive scans by minimizing alignment errors. These have been enhanced by variants like GICP [7], which boost robustness and efficiency. Advanced methods such as CT-ICP [16], KISS-ICP [2], DLO [4] and DLIO [17] all perform scan-to-local-map (or scan-to-submap) registration, leveraging accumulated map information for improved stability. To reduce computational cost, feature-based methods extract geometric structures rather than using entire point clouds. LOAM [3] pioneered this approach by detecting edges and planar features, later extended by LeGO-LOAM[5], F-LOAM [6], and A-LOAM [18], while M-LOAM [33] adapts the framework to multi-LiDAR configurations. In contrast, SuMa++ [19] abandons sparse features in favor of a dense surfel map, each surfel encoding position, normal, temporal stability, and semantic information, for projective ICP registration and dynamic-object filtering.
@@
Despite its advantages, LiDAR-only odometry faces several limitations. In featureless environments, such as tunnels, point cloud registration becomes unreliable due to insufficient constraints. Additionally, LiDAR’s low frame rate makes it less effective for capturing rapid motion dynamics. To mitigate these issues, sensor fusion with complementary modalities is employed to enhance robustness and computational efficiency. LiDAR-Inertial odometry is an approach commonly used as IMU provide high-frequency inertial measurements capturing short-term motion dynamics with minimal latency.
Sensor fusion methods can be categorized depending on how each sensor contributes to the estimation of the pose. In loosely coupled approaches, LiDAR and IMU measurements are processed independently, and their outputs are fused at a later stage. These methods offer modularity and robustness, but may not fully exploit the complementary nature of both sensors. Examples include LOAM [3] or Lego-LOAM [5].
On the other hand, tightly coupled approaches directly integrate raw IMU and LiDAR measurements within the same optimization framework, enabling continuous state estimation and higher accuracy. Methods such as LIO-SAM [1], FastLIO2 [23] and Faster-LIO [24] perform well in feature-sparse environments, where IMU constraints help maintain reliable pose estimates. Despite advances in LiDAR-based odometry, several challenges still hinder its performance. One of the main issues is the high computational cost of processing dense point clouds, making it difficult to execute registration and optimization algorithms efficiently on hardware-constrained platforms. Another significant challenge is the difficulty of extracting discriminative features in homogeneous environments or repetitive geometries, potentially leading to errors in motion estimation. In such scenarios, direct odometry methods, that leverage the full point cloud information without relying on feature extraction, are often considered as an alternative [4] [19]. Furthermore, filtering and data reduction strategies must balance the need to decrease computational load while preserving essential information for localization. Recently, neural implicit representations have been explored as an alternative to traditional geometric or TDF-based maps. Approaches such as NeRF-LOAM [25], PIN-SLAM [26], and KN-LIO [27] jointly optimize odometry and dense mapping within continuous neural fields. These methods leverage signed distance functions or neural point-based representations to achieve accurate pose estimation together with high-fidelity 3D reconstructions, often integrating LiDAR and IMU measurements. Although promising, they typically involve high computational demands and GPU acceleration, which may limit their applicability in resource-constrained robotic platforms.
@@ II-B Distance Fields Computation
Volumetric mapping and SLAM methods primarily utilize discrete voxel grids, hierarchical data structures, probabilistic models, and more recently implicit neural fields.
Voxel-based methods are extensively used due to their simplicity and ease of implementation. Voxblox [28] efficiently generates Euclidean Signed Distance Fields (ESDF) from Truncated Signed Distance Fields (TSDF), establishing a common baseline in volumetric mapping. Additionally, Voxfield [8] refines accuracy through non-projective TSDF fusion complemented by normal estimation from input point-cloud data. Notably, both Voxblox and Voxfield use voxel hashing, which enables sparse spatial memory allocation and helps mitigate scalability issues in large-scale environments. However, these approaches can still suffer from discretization artifacts and access-time variability, particularly in CPU-only systems or high-frequency mapping tasks.
Hierarchical data structures such as OpenVDB [9] have been proposed to mitigate scalability constraints. Approaches leveraging OpenVDB include VDBFusion [29], which integrates TSDF data into OpenVDB. While these hierarchical methods improve scalability and memory usage, their discrete nature still limits smooth interpolation capabilities and continuous optimization.
Implicit neural representations have emerged as continuous and differentiable alternatives to discrete volumetric maps. DeepSDF [30] introduced neural SDFs for continuous surface modeling, and real-time incremental updates were first achieved by iSDF [11]. More recently, HIO-SDF [12] combines global voxel structures with local LiDAR data for online training, while SHINE-Mapping [31] employs a sparse-octree neural SDF for large-scale reconstruction, but it targets mapping rather than online LiDAR–inertial odometry. NeRF-SLAM [32] delivers real-time dense monocular SLAM with hierarchical NeRFs, yet none of them unifies full 6-DoF LiDAR–inertial odometry with on-the-fly distance field mapping on a CPU. Despite these advances, neural methods typically require significant pre-training, heavy computation, and are limited to small-scale SLAM due to memory and inference constraints.
Gaussian Processes (GP) have been proposed as probabilistic distance‐field representations with built-in uncertainty estimation. GP-based mapping methods remain computationally expensive due to covariance matrix inversion, even when hierarchical structures such as Octrees are used. VDB-GPDF [13] embeds logarithmic GP models in OpenVDB to produce continuous ESDFs with uncertainty, but GP inference still limits real-time performance.
@@
In contrast, the approach proposed in this paper, Direct LiDAR-Inertial Odometry (D-LIO), explicitly targets these limitations. D-LIO simplifies the odometry pipeline by eliminating the necessity for explicit LiDAR feature selection and tracking. Moreover, it maintains the environment map update process in constant time, irrespective of map size, significantly enhancing scalability. By directly fusing raw LiDAR point-cloud data into an optimized Fast Truncated Distance Field (Fast-TDF) representation, our method eliminates the additional computational overhead typically required to convert TSDF to ESDF, thus facilitating real-time integration into SLAM systems.
@@ III Fast Truncated Distance Field
Distance Fields (DFs) are able to provide the distance to the closest obstacle from any point in space, no matter how far we are from the obstacle. On the other hand, Truncated Distance Fields (TDFs) are only useful when the coordinate has an obstacle under its truncation distance. While this might be seen as a great advantage of DFs for localization, practical aspects limits and even dismiss such advantage because far points are normally rejected by the outlier rejection process during point cloud registration. Thus, if the truncation distance is properly tuned according to the localization approach at hand, TDFs might be a very efficient representation of the distance field.
@@ III-A Binary kernel
As an illustrative example, let’s assume we have three different distance values represented as a bit-set mask of 8 bits, these values are $00111111$ ($L_{1}=6$), $00001111$ ($L_{1}=4$) and $00000001$ ($L_{1}=1$). The shortest distance among these values can be easily estimated by computing the bitwise-AND operation without performing value comparison:
@@ III-B TDF grid map
We store the TDF map with two levels of hashing: the first level splits space into $1\,\text{m}^{3}$ blocks, and the second level manages the high-resolution cells inside a block. Only blocks that actually receive points are expanded to high resolution (allocated on demand by the second level). Each allocated fine cell stores the $L_{1}$ distance to the closest point of the 3D map to that cell, or the truncation distance if the closest point is far. The truncation distance will be defined by three factors: the number of bits used to represent the bit-set mask, the size of the kernel, and the grid resolution. The number of bits of the mask must be equal to the maximum $L_{1}$ distance into the kernel in order to make the approach efficient, so the truncation distance will be finally given by the size of the kernel and the resolution.
III-C Grid map initialization and access
As previously introduced, we organize the map with a two-level (double) hashing scheme designed for memory efficiency. The first level partitions space into fixed $1\,\text{m}^{3}$ blocks. Many of these blocks will never receive measurements in typical scenes, so they remain lightweight: each block simply holds a pointer that is null until needed. When a point falls inside a block for the first time, that pointer is set to a newly allocated high-resolution container for that block. At the second level, the fractional part of the coordinates (scaled by the chosen resolution) determines the fine cell inside the block where the binary mask value is stored. This on-demand allocation ensures that only blocks receiving points are expanded to high resolution, while all other regions remain unallocated. If the maximum number of fine containers is reached, the oldest one is recycled through a circular buffer, keeping memory bounded and efficient.
To ensure correct operation of the binary kernel with bitwise-AND, high-resolution containers are initialized to the maximum truncation distance, that is, a mask with all bits set. This value will be updated by the binary kernel as soon as the value to store into the grid cell is smaller than the truncation distance. Containers that remain unused are not allocated, and any recycled container (when hitting the memory cap) is re-initialized to the maximum truncation distance before reassignment.
@@ III-E Distance interpolation
Following [34], we make use of trilinear interpolation in order to estimate the distance field beyond the grid cell resolution. However, here, the interpolation is computed online.
@@ IV Direct 3D LiDAR-Inertial Odometry
Figure 2 shows a block diagram of the D-LIO approach. The algorithm incrementally estimates the system pose using an Extended Kalman Filter (EKF) by integrating inertial measurements for pose prediction and point cloud to map registration for pose update. Thus, for each incoming point cloud, the system performs a cloud registration by aligning it with the environment mapped so far based on the Truncated Distance Field (TDF), using the EKF estimate as initial guess. This alignment is used to constrain the EKF’s pose and velocity by means of state update. Whenever the estimated pose variation exceeds a predefined threshold, the TDF map is updated by incorporating new information. This updated map then facilitates a more precise alignment of subsequent point clouds during the registration process.
@@ IV-A Inertial State Integration
Following LiDAR odometry common practices [35][23], inertial sensors can be used to improve state estimation between scans. Inertial integration provides a good initial solution for LiDAR scan alignment, which benefits the subsequent scan-matching process. In addition, high-frequency inertial integration enables LiDAR scan deskewing, which also positively impacts the quality of the scan-matching process. We make use of an EKF with the following state vector:
@@ IV-D Key-framing and map updating
As noted earlier, in case all high-resolution containers are already in use when new points arrive (e.g., as the vehicle enters previously unseen areas), the system removes the information from the oldest $1\,\text{m}^{3}$ block and reassigns its high-resolution container to the newly observed block in the current scan, keeping memory bounded while prioritizing currently visible regions.
@@ V Experimental results
The algorithm has been tested on three datasets, the VIRAL Dataset [37], the Newer College Dataset [38] and the VBR Dataset [39], to evaluate its ability to correctly localize in different environments using diverse trajectories that challenge its performance.
The accurate estimation of odometry has been primarily validated with the VIRAL Dataset, as it features more challenging environments and complex aerial trajectories that test the algorithm’s stability. We evaluate the sequences “eee”,“nya”, and “sbs”, using a DJI M600 Hexacopter equipped with two LiDARs (Ouster OS1-16 gen 1, horizontal and vertical), providing a combined density of approximately 20k points/scan. For the Newer College Dataset (“Quad-Easy” trajectory), we evaluate the reconstruction quality obtained from the TDF map, leveraging existing benchmarks on this dataset that report results across different approaches. Finally, the VBR dataset is used to demonstrate the algorithm’s scalability in a significantly larger environment, showcasing its ability to handle increasing map size and dense point clouds. In this case, we evaluate two trajectories: “diag”, an indoor/outdoor sequence of 1.4 km across three building floors with several loops, and “pincio”, an outdoor 2.5 km trajectory covering a wide open area with dense vegetation. In both datasets, the platform is handheld with an Ouster OS-0-128 LiDAR; with 1024 columns for the Newer College ($\sim$130k points/scan) and 2048 columns for the VBR ($\sim$262k points/scan).
Our results are presented in Tables I and II together with other approaches, such as A-LOAM [18], M-LOAM [33], FAST-LIO2 [23], KISS-ICP [2], LIO-SAM[1] and PIN-SLAM [26]. In all three datasets, the Absolute Translation Error (ATE) was computed using the benchmarking tools provided by the VIRAL dataset to ensure a consistent and reliable trajectory evaluation. For the VBR dataset, we also report the Relative Pose Error (RPE), as it is especially informative for long trajectories. The plots of all trajectories, together with complementary information and 3D reconstructions, are available in the supplementary video [^1] https://www.youtube.com/watch?v=HmrA9YOCZ9w and repository. [^2] https://github.com/robotics-upo/D-LIO.git The experiments were performed on an HP Victus 16 laptop equipped with 32 GB RAM and a 13th generation Intel Core i7-13700H processor. No GPU was used for the experimentation.
V-A Trajectory evaluation
These experimental results made use of Fast-TDF using 64 bit mask encoding together with a grid resolution of 0.05 m, the truncation distance of this kernel is approximately 3 meters. The overall map extent is dataset-dependent; however, by using the proposed coarse layer of $1\,\text{m}^{3}$ blocks, we can scale the map footprint to very large areas at low resolution while allocating high-resolution cells only where points are observed. In our largest runs, we scaled the map to 500×500×100 m. Finally, the rotation/translation thresholds used to update the grid were set to $\lambda=2.0$, $t_{\text{th}}=1\,\mathrm{m}$, and $q_{\text{th}}=\pi\,\mathrm{rad}$. As shown in Table I, our method attains the lowest ATE in most sequences and ranks second-best in the remainder, highlighting its consistent accuracy and robustness across diverse and challenging trajectories.
Please, notice that LIO-SAM and PIN-SLAM performs back-end optimization, which makes the comparison with pure odometry not entirely fair. However, we have included its results to demonstrate that our approach, despite not using back-end optimization, achieves comparable performance.